# Exploratory Randomization for Discrete-Time Linear Exponential Quadratic Gaussian (LEQG) Problem

### v1

@@ 4.1 Analysis of the sufficiency conditions for the existence of a saddle point
Given scalar model parameters $A,B,M>0,\theta>0$ and choosing for simplicity of presentation $Q=m=n=0$, an explicit sufficient condition for the stationary point $(u^{*},\eta^{*},\gamma*)$ from Proposition 3.3 to be a saddle point for $V_{t}(x_{t})$ with a $\min$ in $u^{*}$ and a joint $\max$ in $\nu^{*}=(\gamma^{*},\eta^{*})$ is given by the following three requirements:
@@
    $\Xi_{t}^{-1}>\max\left[2\theta\,N_{t}+B^{2}P_{t+1},2\theta\,N_{t}+\frac{\Lambda_{t}^{-1}B^{2}P_{t+1}}{P_{t+1}-\Lambda_{t}^{-1}}\right]=\left\{\begin{array}[]{lcl}2\theta\,N_{t}+B^{2}P_{t+1}&\mbox{if}&P_{t+1}>0\\ 2\theta\,N_{t}+\frac{\Lambda_{t}^{-1}B^{2}P_{t+1}}{P_{t+1}-\Lambda_{t}^{-1}}&\mbox{if}&P_{t+1}<0\end{array}\right.$
@@
Procedure: Given is a discrete time period $t=0,\cdots,T$
@@
    Compute $P_{T-1},p_{T-1},r_{T-1}$ according to (i), (ii), eqiii in Theorem 3.5 and repeat steps 4. and 5. successively for the times $T-1,T-2,\cdots,1,0$.
@@
    Run the system from $t=0$ to $t=T$ with the control $u^{*}$ given by (3.3) of Proposition 3.3 and with an unbiased randomization noise $v_{t}$ given by a zero-mean Gaussian with variance $\Xi_{t}$. For a simulation, choose a system noise $w_{t}\sim{\cal N}(\gamma_{t}^{*},\Lambda_{t})$.
@@
Since the cost function $I(u;T,\theta)$ is positive, the value function $V_{t}(x_{t})$ will also be positive, so the most likely sign for $P_{t}$ is positive. By i) in Proposition 4.1, the control penalization coefficient $N_{t}$ can then be chosen arbitrarily as a positive matrix.
    If we run a real system, the system noise $w_{t}$ might have a covariance $\Lambda_{t}$ with $\Lambda_{t}^{-1}<P_{t+1}$. In such a case we cannot be sure that our controls $u^{*}$ and $\gamma^{*}$ lead to a saddle point as required in (3.12) (see also (3.3)). On the other hand, if we simulate the system, the system noise has to have a variance that, in the scalar case, is smaller for larger values of $P_{t}$. Furthermore, the variance $\Xi_{t}$ of the unbiased randomization noise has to be such that
@@
    Since our optimal controls are parametrized by the system parameters $A,a,B$, we may also run a policy gradient method and update the controls directly without previously estimating the system parameters sequentially.
@@ 4.2 Conclusions
We have seen that a LEQG problem with a randomized control emulating exploration can, via the energy-entropy duality, be reduced to a risk-neutral LQG problem with an additive penalization. This LQG problem takes the form of a stochastic game where one player is the original controller of the LEQG problem. The other two players result from the randomization and the duality relation. They appear due to the reduction of the problem from LEQG to LQG. We have also investigated the conditions required to have an LQG game problem. While the data for the model and the problem setup can be chosen freely, the state noise and the randomization have to satisfy certain conditions that, at the generic time $t$, involve the previously computed coefficient $P_{t+1}$ in the leading term of the quadratic value function $V_{t+1}(x_{t+1})$. This is one of the reasons why we considered the system and randomization noises $\Lambda_{t}$ and $\Xi_{t}$ as time-dependent. The coefficient $P_{t}$ is deterministic and can be determined a priori via the backward recursion (i) in Theorem 3.5 based on the estimated model data. Hence, the above-mentioned conditions can actually be checked when, given the other data, one chooses/verifies the state noise and the randomization intensity. Essentially, the state noise should not be too diffuse if $P_{t}$ is large. At the same time, the randomization should have no bias and not be more diffuse than a bound given by $P_{t}$, the risk sensitivity parameter $\theta$ and the control weight factor $N_{t}$.

### v2

@@ 4.1 Analysis of the sufficiency conditions for the existence of a saddle point
Given scalar model parameters $A>0,B>0,M>0,\theta>0$ and choosing for simplicity of presentation $Q=m=n=0$, an explicit sufficient condition for the stationary point $(u^{*},\eta^{*},\gamma*)$ from Proposition 3.3 to be a saddle point for $V_{t}(x_{t})$ with a $\min$ in $u^{*}$ and a joint $\max$ in $\nu^{*}=(\gamma^{*},\eta^{*})$ is given by the following three requirements:
@@
    $\Xi_{t}^{-1}>\max\left[2\theta\,N_{t}+B^{2}P_{t+1},2\theta\,N_{t}+\frac{\Lambda_{t}^{-1}B^{2}P_{t+1}}{P_{t+1}-\Lambda_{t}^{-1}}\right]=\begin{cases}2\theta\,N_{t}+B^{2}P_{t+1}&\text{if }P_{t+1}>0\\ 2\theta\,N_{t}+\frac{\Lambda_{t}^{-1}B^{2}P_{t+1}}{P_{t+1}-\Lambda_{t}^{-1}}&\text{if }P_{t+1}<0\end{cases}$
@@
Procedure ( recursion): Given is a discrete time period $t=0,\cdots,T$
@@
    Compute $P_{T-1},p_{T-1},r_{T-1}$ according to (3.43), (3.44), (3.45) in Theorem 3.5 and repeat steps 4. and 5. successively for the times $T-1,T-2,\cdots,1,0$.
@@
    Run the system from $t=0$ to $t=T$ with the control $u^{*}$ given by (3.23) of Proposition 3.3 and with an unbiased randomization noise $v_{t}$ given by a zero-mean Gaussian with variance $\Xi_{t}$. For a simulation, choose the most recent estimated values of A and B and a system noise $w_{t}\sim\mathcal{N}(\gamma_{t}^{*},\Lambda_{t})$.
@@
Since the criterion $I(u;T,\theta)$ is positive, the value function $V_{t}(x_{t})$ will also be positive, so the most likely sign for $P_{t}$ is positive. By i) in Proposition 4.1, the control penalization coefficient $N_{t}$ can then be chosen arbitrarily as a positive matrix. If we run a real system, the system noise $w_{t}$ might have a covariance $\Lambda_{t}$ with $\Lambda_{t}^{-1}<P_{t+1}$. In such a case, we cannot be sure that our controls $u^{*}$ and $\gamma^{*}$ lead to a saddle point as required in (3.12) (see also (3.22)). On the other hand, if we simulate the system, the system noise has to have a variance that, in the scalar case, is smaller for larger values of $P_{t}$. Furthermore, the variance $\Xi_{t}$ of the unbiased randomization noise has to be such that
@@
4.2 Numerical Example
We provide here a simple numerical example for the scalar case considered in the previous section 4.1. In addition to providing illustrative numbers, one major interest of this example is that it shows that the setup of Proposition 4.1 is sufficient, but not necessary.
Table 2 reports the model parameters, which are all constant. This choice of parameters departs from the setup of Proposition 4.1 in two ways. First, A=0.2<0A=-0.2<0, implying that $x_{t}$ has a tendency to revert to its long-term mean. Second, $Q=1\neq 0$ to model an interaction between the state and control in the cost function [^2] This model specification is consistent with the continuous-time risk-sensitive models arising in investment management (see for instance Kuroda and Nagai, 2002; Davis and Lleo, 2014; Lleo and Runggaldier, 2024). Nevertheless, we will see below that Assumption 3.2 is satisfied, and so, by Proposition 3.7, the game has a unique saddle point at $(u^{*},\gamma^{*},\eta^{*})$.
Global Parameters System Dynamics Cost Function
Parameter Value Parameter Value Parameter Value
$\theta$ 1 $a$ 0 $M$ 2
$T$ 25 AA -0.2 NN 2
$B$ 0.4 $Q$ 1
$\Lambda$ 0.15 mm 0
$\Xi$ 0.15 $n$ 0
$x_{0}$ 1 $M_{T}$ 4
$m_{T}$ 0
Table 2: Model parameters
The solution of the game is displayed in Table 3 below, which shows the evolution of the state variable $x_{t}$ under both the $\mathbb{P}$ and $\mathbb{P}^{\gamma^{*},\eta^{*}}$ measures, of the value function $V_{t}(x_{t})$ and its parameters $P_{t},p_{t},r_{t}$, of the optimal controls $u^{*}_{t}$, $\gamma^{*}_{t}$, and $\eta^{*}_{t}$, and of the intermediate Fraktur parameters required to check Assumption 3.2. As $\mathfrak{B}^{(1)}_{t}>0,\;t=1,\ldots,T$, Assumption 3.2(i) is satisfied. Next, we check that Assumption 3.2(ii), that is,
$H^{\bar{\gamma},\bar{\eta}}_{t}=-\begin{pmatrix}\mathfrak{B}^{(2)}_{t+1}&\mathfrak{C}_{t+1}^{\prime}\\ \mathfrak{C}_{t+1}&\mathfrak{B}^{(3)}_{t+1}\end{pmatrix}>0,$ (4.1)
is also satisfied. To check this assumption, we use Sylvester’s criterion. From Table 3, we have $\mathfrak{B}^{(2)}_{t}<0,\;t=0,\ldots,T$. We also see that the determinant $\det\left(-H^{\bar{\gamma},\bar{\eta}}_{t}\right)=\mathfrak{B}^{(2)}_{t}\mathfrak{B}^{(3)}_{t}-\mathfrak{C}_{t}^{2}$, is strictly positive for $t=0,\ldots,T$. Therefore, Assumption 3.2 is fully satisfied and the game has a unique saddle point at $(u^{*}_{t},\gamma^{*}_{t},\eta^{*}_{t})$. Notice also that, despite the different setups, the sufficient conditions i) to iii) of Proposition 4.1 would be satisfied for this example.
$t$ $x_{t}$ under $\mathbb{P}$ $x_{t}$ under $\mathbb{P}^{\gamma^{*},\eta^{*}}$ $P_{t}$ $p_{t}$ $r_{t}$ $V_{t}(x_{t})$ $u^{*}_{t}$ $\gamma^{*}_{t}$ $\eta^{*}_{t}$ $\mathfrak{B}^{(1)}_{t}$ $\mathfrak{B}^{(2)}_{t}$ $\mathfrak{B}^{(3)}_{t}$ $\mathfrak{C}_{t}$ $\det\left(-H^{\bar{\gamma},\bar{\eta}}_{t}\right)$
0 1.0000 1.0000 4.5811 0.0000 17.3880 19.6786 0.0270 -0.4156 -0.0000
1 -0.2655 -0.6782 4.5811 0.0000 16.6894 17.7430 -0.0183 0.2818 0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
2 -0.1205 0.2024 4.5811 0.0000 15.9909 16.0847 0.0055 -0.0841 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
3 -0.0357 -0.1270 4.5811 0.0000 15.2923 15.3293 -0.0034 0.0528 0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
4 -0.0629 0.0018 4.5811 0.0000 14.5938 14.5938 0.0000 -0.0007 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
5 -0.0581 -0.1008 4.5811 0.0000 13.8952 13.9185 -0.0027 0.0419 0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
6 0.0225 0.1376 4.5811 0.0000 13.1967 13.2400 0.0037 -0.0572 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
7 0.3837 0.2704 4.5811 0.0000 12.4981 12.6655 0.0073 -0.1123 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
8 0.2583 0.1856 4.5811 0.0000 11.7995 11.8785 0.0050 -0.0771 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
9 -0.3427 -0.3417 4.5811 0.0000 11.1010 11.3685 -0.0092 0.1420 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
10 0.0247 0.2576 4.5811 0.0000 10.4024 10.5544 0.0070 -0.1070 0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
11 0.0176 -0.1256 4.5811 0.0000 9.7039 9.7400 -0.0034 0.0522 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
12 0.0367 0.0863 4.5811 0.0000 9.0053 9.0224 0.0023 -0.0359 -0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
13 0.0111 -0.1089 4.5811 0.0000 8.3067 8.3339 -0.0029 0.0453 0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
14 0.1961 0.3188 4.5811 0.0000 7.6082 7.8409 0.0086 -0.1325 0.0000 4.7330 -2.0855 -1.9337 1.8324 0.6749
15 -0.0315 -0.1993 4.5811 0.0000 6.9096 7.0006 -0.0054 0.0828 -0.0000 4.7330 -2.0856 -1.9337 1.8324 0.6750
16 -0.1065 0.0364 4.5811 0.0000 6.2111 6.2141 0.0010 -0.0151 0.0000 4.7330 -2.0856 -1.9337 1.8324 0.6751
17 0.1562 0.1640 4.5810 0.0000 5.5125 5.5741 0.0044 -0.0681 -0.0000 4.7330 -2.0857 -1.9337 1.8324 0.6755
18 -0.0318 -0.1949 4.5807 0.0000 4.8140 4.9010 -0.0052 0.0809 -0.0000 4.7329 -2.0860 -1.9338 1.8323 0.6765
19 -0.0052 0.0639 4.5800 0.0000 4.1156 4.1249 0.0017 -0.0265 0.0000 4.7328 -2.0867 -1.9339 1.8320 0.6792
20 -0.0517 -0.0946 4.5780 0.0000 3.4173 3.4378 -0.0025 0.0392 0.0000 4.7325 -2.0887 -1.9342 1.8312 0.6867
21 0.1290 0.2099 4.5725 0.0000 2.7195 2.8202 0.0051 -0.0863 0.0000 4.7316 -2.0942 -1.9351 1.8290 0.7073
22 0.0813 -0.0077 4.5573 0.0000 2.0230 2.0231 -0.0001 0.0031 -0.0000 4.7292 -2.1094 -1.9375 1.8229 0.7638
23 0.2333 0.2933 4.5150 0.0000 1.3302 1.5244 0.0015 -0.1122 -0.0000 4.7224 -2.1516 -1.9443 1.8060 0.9217
24 0.0781 -0.0031 4.3929 0.0000 0.6480 0.6480 0.0001 0.0010 0.0000 4.7029 -2.2738 -1.9638 1.7571 1.3778
25 0.2215 0.2497 4.0000 0.0000 0.0000 0.1247 4.6400 -2.6667 -2.0267 1.6000 2.8444
Table 3: One run of the numerical example
Remark 9.
As expected, in Table 3, $\eta^{*}$ is practically 0.
4.3 Integration with Policy Gradient and Actor–Critic
In Sections 2 and 3, we introduced the LEQG control problem, expressed it as an equivalent LQG game using the duality between free energy and entropy, and derived the solution to this game. Section 4.1 then analyzed sufficient conditions for the game to have a unique saddle point and proposed a procedure to estimate sequentially the model parameters and check that the sufficient requirements of Proposition 4.1 are satisfied.
As an alternative, we could also have developed and implemented a policy gradient approach, such as an actor-critic algorithm, to solve the problem numerically using only realized sample trajectories of the state space $x_{t}$ and cost $G_{T}$. A full and rigorous treatment is beyond the scope of this paper and should be the focus of further research. We thus only sketch the main ideas here.
Policy gradient methods have gained popularity in recent years following groundbreaking advances in the field of reinforcement learning and a string of highly successful applications, most strikingly to the game of Go. We refer the reader to Chapter 13 in the classic text by Sutton and Barto (2018).
All policy gradient methods follow the same essential steps. First, approximate the policy using a conveniently parametrized function. Second, compute the gradient of a given performance measure, usually a value function, with respect to the parameters of the function approximating the controls. Third, update the parameters by performing a gradient descent update if the aim is to minimize the performance measure, or a gradient ascent update if the aim is to maximize the performance measure.
Actor-critic methods build on this approach to update both the policy and the value function. The actor performs the three policy gradient steps to update the parameters of the policy approximation, at the current estimate for the value function, while the critic performs three similar steps to update the parameters of an approximation of the value function, at the current estimate for the policy.
To implement a simple policy gradient method in our context, a natural approach is to extend the approach by Hambly et al. (2021, 2023) on policy gradient methods for linear quadratic regulators and linear-quadratic games. Their work captures the essential features of LQG control problems and games, namely the quadratic cost in the state and in the controls, and the linear dynamics of the state variable, and provides strong convergence results. However, it does not consider some additional features that are present in our model and are important for applications, in particular to finance, such as the interaction between the state and control in the cost function, first-order terms in the cost function, and the possibility of a mean reversion to a long-term mean different from 0 for the state. Hence, further work is necessary.
For our LQG game, the optimal controls are affine in the state, and the value function is quadratic. We can therefore use the following parametrization [^3] Note that this is a full parametrization. We could use the fact that $\eta^{*}=0$ to substantially simplify the mathematical exposition and reduce the amount of numerical work.:
$u_{t}=\mathtt{D}_{t}x_{t}+\mathtt{d}_{t}\qquad\bar{\gamma}_{t}=\mathtt{E}_{t}x_{t}+\mathtt{e}_{t}\qquad\bar{\eta}_{t}=\mathtt{F}_{t}x_{t}+\mathtt{f}_{t}$ (4.2)
for $t=0,\ldots,T-1$, where $\mathtt{D}_{t}\in\mathbb{R}^{d_{u}\times d_{x}}$, $\mathtt{d}_{t}\in\mathbb{R}^{d_{u}}$, $\mathtt{E}_{t}\in\mathbb{R}^{d_{x}\times d_{x}}$, $\mathtt{e}_{t}\in\mathbb{R}^{d_{x}}$, $\mathtt{F}_{t}\in\mathbb{R}^{d_{u}\times d_{x}}$, $\mathtt{f}_{t}\in\mathbb{R}^{d_{u}}$. Define $\mathtt{\mathbf{D}}:=\left(\mathtt{D}_{0},\ldots,\mathtt{D}_{T-1}\right)$ with similar definitions for $\mathtt{\mathbf{d}},\mathtt{\mathbf{E}},\mathtt{\mathbf{e}},\mathtt{\mathbf{F}},\mathtt{\mathbf{d}}$, and let $\mathbf{K}=\left(\mathtt{\mathbf{D}},\mathtt{\mathbf{d}},\mathtt{\mathbf{E}},\mathtt{\mathbf{e}},\mathtt{\mathbf{F}},\mathtt{\mathbf{f}}\right)$. Furthermore, we take the critic as
$V_{t}(x)=x_{t}^{\prime}\mathtt{P}_{t}x+x_{t}^{\prime}\mathtt{p}_{t}+\mathtt{r}_{t},$ (4.3)
where $\mathtt{P}_{t}\in\mathbb{R}^{d_{x}\times d_{x}},\mathtt{p}_{t}\in\mathbb{R}^{d_{x}}$, and $\mathtt{r}_{t}\in\mathbb{R}$.
Based on (3.12), the objective function $C$, as a function of the policy parameters, is
$C(\mathtt{D},\mathtt{d},\mathtt{E},\mathtt{e},\mathtt{F},\mathtt{f})$ (4.4)
[…]
$\left.\left.+\left(\mathtt{F}_{t}x_{t}+\mathtt{f}_{t}\right)^{\prime}\Xi_{t}^{-1}\left(\mathtt{F}_{t}x_{t}+\mathtt{f}_{t}\right)\right\}\right]$
With this notation, Hambly et al.’s so-called Natural Policy Gradient updating rules (extended to include a zeroth order term) for a sequence of episodes $m=1,\ldots,M$ are:
$\mathbf{K}^{(m+1)}=\mathbf{K}^{(m)}\pm\delta_{m}\,\widehat{\nabla}_{\mathbf{K}}C\!\left(\mathbf{K}^{(m)}\right)\,\left(\widehat{\Sigma}^{(m)}_{x}\right)^{-1},$ (4.5)
for $t=0,1,\ldots,T-1$ and where $\delta$ is the step size. For the case of unknown model parameters, $\widehat{\nabla}_{\mathbf{K}}C$ is the empirical policy gradient, and $\Sigma^{\mathbf{K}}_{t}:=\mathbf{E}\left[x^{\mathbf{K}}_{t}\left(x^{\mathbf{K}}_{t}\right)^{\prime}\right]$ is the empirical state covariance (serving as a natural preconditioner). The sign is negative (descent) for the minimizing controller parameters $(\mathtt{D},\mathtt{d})$ and positive (ascent) for the adversarial parameters $(\mathtt{E},\mathtt{e},\mathtt{F},\mathtt{f})$.
From there, the next step is to consider a full actor-critic approach, where the gradient update to the policy approximation combines with a gradient update to the value function approximation. The critic parameters $(\mathtt{P}_{t},\mathtt{p}_{t},\mathtt{r}_{t})$ are updated on a faster timescale by Temporal Differencing. The key here is to establish that the joint gradient updates will converge to the optimal controls and the optimal value function.
Finally, we note that global convergence results exist for policy-gradient methods for LQR problems– see (Fazel et al., 2018; Hambly et al., 2021) for control problems and Hambly et al. (2023) for linear-quadratic games. On the actor-critic front, Alacaoglu et al. (2022) obtained promising results for 2-player Markov games.
@@ 4.4 Conclusions
We have seen that an LEQG problem with a randomized control emulating exploration can, via the energy-entropy duality, be reduced to a risk-neutral LQG problem with an additive penalization. This LQG problem takes the form of a stochastic game where one player is the original controller of the LEQG problem. The other two players result from the duality relation. They appear due to the reduction of the problem from LEQG to LQG. We have also investigated sufficient conditions to have an LQG game problem. While in the recursive procedure the data for the model and the problem setup can be chosen freely, the state noise and the randomization should satisfy certain conditions that, at the generic time $t$, involve the previously computed coefficient $P_{t+1}$ in the leading term of the quadratic value function $V_{t+1}(x_{t+1})$. This is one of the reasons why we considered the system and randomization noises $\Lambda_{t}$ and $\Xi_{t}$ as time-dependent. The coefficient $P_{t}$ is deterministic and can be determined via the backward recursion (3.43) in Theorem 3.5 based on the estimated model data. Hence, the above-mentioned conditions can actually be checked when, given the other data, one chooses/verifies the state noise and the randomization intensity. Essentially, the state noise should not be too diffuse if $P_{t}$ is large. At the same time, the randomization should have no bias and not be more diffuse than a bound given by $P_{t}$, the risk sensitivity parameter $\theta$, and the control weight factor $N_{t}$.